OpenStudy (anonymous):

Use the function to answer the question. f(x) = 4x - 8 what is the value of f^-1(-2)? please explain

OpenStudy (one098):

So plug in, (-2) into the x's in the function above. Multiply 4 times -2 and you get.. then you subtract -8 and 8 and the answer is...?

OpenStudy (anonymous):

-16??

OpenStudy (one098):

Exactly!

OpenStudy (anonymous):

my teacher says thats wrong though

OpenStudy (mathmath333):

first find f inverse \(\large\tt \color{black}{f(x) = 4x - 8}\) \(\large\tt \color{black}{y = 4x - 8}\) \(\large\tt \color{black}{y +8= 4x}\) \(\large\tt \color{black}{\dfrac{y +8}{4}= x}\) swap x with y \(\large\tt \color{black}{\dfrac{x+8}{4}= y}\) \(\large\tt \color{black}{f^{-1}(x) = \dfrac{x+8}{4}}\) \(\large\tt \color{black}{f^{-1}(-2) = \dfrac{-2+8}{4}}\) \(\large\tt \color{black}{f^{-1}(-2) = \dfrac{6}{4}}\) \(\large\tt \color{black}{f^{-1}(-2) = \dfrac{3}{2}}\)

OpenStudy (anonymous):

thank you @mathmath333

OpenStudy (mathmath333):

welcome